\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Seiberg:1994pq}
N.~Seiberg, ``{Electric - magnetic duality in supersymmetric nonAbelian gauge
  theories},'' \href{http://dx.doi.org/10.1016/0550-3213(94)00023-8}{{\em Nucl.
  Phys.} {\bfseries B435} (1995) 129--146},
\href{http://arxiv.org/abs/hep-th/9411149}{{\ttfamily arXiv:hep-th/9411149
  [hep-th]}}.
%%CITATION = HEP-TH/9411149;%%.

\bibitem{Smilga:1994tb}
A.~V. Smilga and J.~J.~M. Verbaarschot, ``{Spectral sum rules and finite volume
  partition function in gauge theories with real and pseudoreal fermions},''
  \href{http://dx.doi.org/10.1103/PhysRevD.51.829}{{\em Phys. Rev.} {\bfseries
  D51} (1995) 829--837},
\href{http://arxiv.org/abs/hep-th/9404031}{{\ttfamily arXiv:hep-th/9404031
  [hep-th]}}.
%%CITATION = HEP-TH/9404031;%%.

\bibitem{Kogut:1999iv}
J.~B. Kogut, M.~A. Stephanov, and D.~Toublan, ``{On two color QCD with baryon
  chemical potential},''
  \href{http://dx.doi.org/10.1016/S0370-2693(99)00971-5}{{\em Phys. Lett.}
  {\bfseries B464} (1999) 183--191},
\href{http://arxiv.org/abs/hep-ph/9906346}{{\ttfamily arXiv:hep-ph/9906346
  [hep-ph]}}.
%%CITATION = HEP-PH/9906346;%%.

\bibitem{Magnea:1999iv}
U.~Magnea, ``{The Orthogonal ensemble of random matrices and QCD in
  three-dimensions},'' \href{http://dx.doi.org/10.1103/PhysRevD.61.056005}{{\em
  Phys. Rev. D} {\bfseries 61} (2000) 056005},
  \href{http://arxiv.org/abs/hep-th/9907096}{{\ttfamily arXiv:hep-th/9907096}}.

\bibitem{Dunne:2002vb}
G.~V. Dunne and S.~M. Nishigaki, ``{Two color QCD in 3-D at finite baryon
  density},'' \href{http://dx.doi.org/10.1016/S0550-3213(03)00034-8}{{\em Nucl.
  Phys. B} {\bfseries 654} (2003) 445--465},
  \href{http://arxiv.org/abs/hep-ph/0210219}{{\ttfamily arXiv:hep-ph/0210219}}.

\bibitem{Seiberg:1994bz}
N.~Seiberg, ``{Exact results on the space of vacua of four-dimensional SUSY
  gauge theories},'' \href{http://dx.doi.org/10.1103/PhysRevD.49.6857}{{\em
  Phys. Rev. D} {\bfseries 49} (1994) 6857--6863},
  \href{http://arxiv.org/abs/hep-th/9402044}{{\ttfamily arXiv:hep-th/9402044}}.

\bibitem{Spiridonov:2014cxa}
V.~P. Spiridonov and G.~S. Vartanov, ``{Vanishing superconformal indices and
  the chiral symmetry breaking},''
  \href{http://dx.doi.org/10.1007/JHEP06(2014)062}{{\em JHEP} {\bfseries 06}
  (2014) 062},
\href{http://arxiv.org/abs/1402.2312}{{\ttfamily arXiv:1402.2312 [hep-th]}}.
%%CITATION = ARXIV:1402.2312;%%.

\bibitem{Intriligator:1996ex}
K.~A. Intriligator and N.~Seiberg, ``{Mirror symmetry in three-dimensional
  gauge theories},'' \href{http://dx.doi.org/10.1016/0370-2693(96)01088-X}{{\em
  Phys. Lett.} {\bfseries B387} (1996) 513--519},
\href{http://arxiv.org/abs/hep-th/9607207}{{\ttfamily arXiv:hep-th/9607207
  [hep-th]}}.
%%CITATION = HEP-TH/9607207;%%.

\bibitem{deBoer:1996mp}
J.~de~Boer, K.~Hori, H.~Ooguri, and Y.~Oz, ``{Mirror symmetry in
  three-dimensional gauge theories, quivers and D-branes},''
  \href{http://dx.doi.org/10.1016/S0550-3213(97)00125-9}{{\em Nucl. Phys. B}
  {\bfseries 493} (1997) 101--147},
  \href{http://arxiv.org/abs/hep-th/9611063}{{\ttfamily arXiv:hep-th/9611063}}.

\bibitem{Aharony:1997bx}
O.~Aharony, A.~Hanany, K.~A. Intriligator, N.~Seiberg, and M.~J. Strassler,
  ``{Aspects of N=2 supersymmetric gauge theories in three-dimensions},''
  \href{http://dx.doi.org/10.1016/S0550-3213(97)00323-4}{{\em Nucl. Phys.}
  {\bfseries B499} (1997) 67--99},
\href{http://arxiv.org/abs/hep-th/9703110}{{\ttfamily arXiv:hep-th/9703110
  [hep-th]}}.
%%CITATION = HEP-TH/9703110;%%.

\bibitem{deBoer:1997kr}
J.~de~Boer, K.~Hori, and Y.~Oz, ``{Dynamics of N=2 supersymmetric gauge
  theories in three-dimensions},''
  \href{http://dx.doi.org/10.1016/S0550-3213(97)00328-3}{{\em Nucl. Phys. B}
  {\bfseries 500} (1997) 163--191},
  \href{http://arxiv.org/abs/hep-th/9703100}{{\ttfamily arXiv:hep-th/9703100}}.

\bibitem{Amariti:2016kat}
A.~Amariti, D.~Orlando, and S.~Reffert, ``{String theory and the 4D/3D
  reduction of Seiberg duality. A review},''
  \href{http://dx.doi.org/10.1016/j.physrep.2017.08.002}{{\em Phys. Rept.}
  {\bfseries 705-706} (2017) 1--53},
\href{http://arxiv.org/abs/1611.04883}{{\ttfamily arXiv:1611.04883 [hep-th]}}.
%%CITATION = ARXIV:1611.04883;%%.

\bibitem{Dolan:2011rp}
F.~A.~H. Dolan, V.~P. Spiridonov, and G.~S. Vartanov, ``{From 4d superconformal
  indices to 3d partition functions},''
  \href{http://dx.doi.org/10.1016/j.physletb.2011.09.007}{{\em Phys. Lett.}
  {\bfseries B704} (2011) 234--241},
\href{http://arxiv.org/abs/1104.1787}{{\ttfamily arXiv:1104.1787 [hep-th]}}.
%%CITATION = ARXIV:1104.1787;%%.

\bibitem{Aharony:2013dha}
O.~Aharony, S.~S. Razamat, N.~Seiberg, and B.~Willett, ``{3d dualities from 4d
  dualities},'' \href{http://dx.doi.org/10.1007/JHEP07(2013)149}{{\em JHEP}
  {\bfseries 07} (2013) 149},
\href{http://arxiv.org/abs/1305.3924}{{\ttfamily arXiv:1305.3924 [hep-th]}}.
%%CITATION = ARXIV:1305.3924;%%.

\bibitem{Amariti:2015xna}
A.~Amariti, ``{4d/3d reduction of s-confining theories: the role of the
  \textquotedblleft{}exotic\textquotedblright{} D instantons},''
  \href{http://dx.doi.org/10.1007/JHEP02(2016)139}{{\em JHEP} {\bfseries 02}
  (2016) 139}, \href{http://arxiv.org/abs/1507.05623}{{\ttfamily
  arXiv:1507.05623 [hep-th]}}.

\bibitem{Gahramanov:2015cva}
I.~Gahramanov and V.~P. Spiridonov, ``{The star-triangle relation and 3d
  superconformal indices},''
  \href{http://dx.doi.org/10.1007/JHEP08(2015)040}{{\em JHEP} {\bfseries 08}
  (2015) 040},
\href{http://arxiv.org/abs/1505.00765}{{\ttfamily arXiv:1505.00765 [hep-th]}}.
%%CITATION = ARXIV:1505.00765;%%.

\bibitem{Gahramanov:2016ilb}
I.~Gahramanov and A.~P. Kels, ``{The star-triangle relation, lens partition
  function, and hypergeometric sum/integrals},''
\href{http://arxiv.org/abs/1610.09229}{{\ttfamily arXiv:1610.09229 [math-ph]}}.
%%CITATION = ARXIV:1610.09229;%%.

\bibitem{Kels:2015bda}
A.~P. Kels, ``{New solutions of the star - triangle relation with discrete and
  continuous spin variables},''
  \href{http://dx.doi.org/10.1088/1751-8113/48/43/435201}{{\em J. Phys.}
  {\bfseries A48} no.~43, (2015) 435201},
\href{http://arxiv.org/abs/1504.07074}{{\ttfamily arXiv:1504.07074 [math-ph]}}.
%%CITATION = ARXIV:1504.07074;%%.

\bibitem{Krattenthaler:2011da}
C.~Krattenthaler, V.~P. Spiridonov, and G.~S. Vartanov, ``{Superconformal
  indices of three-dimensional theories related by mirror symmetry},''
  \href{http://dx.doi.org/10.1007/JHEP06(2011)008}{{\em JHEP} {\bfseries 06}
  (2011) 008},
\href{http://arxiv.org/abs/1103.4075}{{\ttfamily arXiv:1103.4075 [hep-th]}}.
%%CITATION = ARXIV:1103.4075;%%.

\bibitem{Gahramanov:2016wxi}
I.~Gahramanov and H.~Rosengren, ``{Basic hypergeometry of supersymmetric
  dualities},'' \href{http://dx.doi.org/10.1016/j.nuclphysb.2016.10.004}{{\em
  Nucl. Phys.} {\bfseries B913} (2016) 747--768},
\href{http://arxiv.org/abs/1606.08185}{{\ttfamily arXiv:1606.08185 [hep-th]}}.
%%CITATION = ARXIV:1606.08185;%%.

\bibitem{Bhattacharya:2008zy}
J.~Bhattacharya, S.~Bhattacharyya, S.~Minwalla, and S.~Raju, ``{Indices for
  Superconformal Field Theories in 3,5 and 6 Dimensions},''
  \href{http://dx.doi.org/10.1088/1126-6708/2008/02/064}{{\em JHEP} {\bfseries
  02} (2008) 064}, \href{http://arxiv.org/abs/0801.1435}{{\ttfamily
  arXiv:0801.1435 [hep-th]}}.

\bibitem{Bhattacharya:2008bja}
J.~Bhattacharya and S.~Minwalla, ``{Superconformal Indices for N = 6 Chern
  Simons Theories},''
  \href{http://dx.doi.org/10.1088/1126-6708/2009/01/014}{{\em JHEP} {\bfseries
  01} (2009) 014}, \href{http://arxiv.org/abs/0806.3251}{{\ttfamily
  arXiv:0806.3251 [hep-th]}}.

\bibitem{Kim:2009wb}
S.~Kim, ``{The Complete superconformal index for N=6 Chern-Simons theory},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2009.06.025}{{\em Nucl. Phys. B}
  {\bfseries 821} (2009) 241--284},
  \href{http://arxiv.org/abs/0903.4172}{{\ttfamily arXiv:0903.4172 [hep-th]}}.
  [Erratum: Nucl.Phys.B 864, 884 (2012)].

\bibitem{Imamura:2011su}
Y.~Imamura and S.~Yokoyama, ``{Index for three dimensional superconformal field
  theories with general R-charge assignments},''
  \href{http://dx.doi.org/10.1007/JHEP04(2011)007}{{\em JHEP} {\bfseries 04}
  (2011) 007}, \href{http://arxiv.org/abs/1101.0557}{{\ttfamily arXiv:1101.0557
  [hep-th]}}.

\bibitem{Willett:2016adv}
B.~Willett, ``{Localization on three-dimensional manifolds},''
  \href{http://dx.doi.org/10.1088/1751-8121/aa612f}{{\em J. Phys.} {\bfseries
  A50} no.~44, (2017) 443006},
\href{http://arxiv.org/abs/1608.02958}{{\ttfamily arXiv:1608.02958 [hep-th]}}.
%%CITATION = ARXIV:1608.02958;%%.

\bibitem{Csaki:2014cwa}
C.~Cs\'aki, M.~Martone, Y.~Shirman, P.~Tanedo, and J.~Terning, ``{Dynamics of
  3D SUSY Gauge Theories with Antisymmetric Matter},''
  \href{http://dx.doi.org/10.1007/JHEP08(2014)141}{{\em JHEP} {\bfseries 08}
  (2014) 141}, \href{http://arxiv.org/abs/1406.6684}{{\ttfamily arXiv:1406.6684
  [hep-th]}}.

\bibitem{Nii:2018erm}
K.~Nii, ``{3d s-confinement for three-index matters},''
  \href{http://dx.doi.org/10.1007/JHEP11(2018)099}{{\em JHEP} {\bfseries 11}
  (2018) 099}, \href{http://arxiv.org/abs/1805.06369}{{\ttfamily
  arXiv:1805.06369 [hep-th]}}.

\bibitem{Nii:2019ebv}
K.~Nii, ``{On s-confinement in 3d $\mathcal{N}=2$ gauge theories with
  anti-symmetric tensors},'' \href{http://arxiv.org/abs/1906.03908}{{\ttfamily
  arXiv:1906.03908 [hep-th]}}.

\bibitem{Amariti:2022wae}
A.~Amariti and S.~Rota, ``{3d N=2 SO/USp adjoint SQCD: s-confinement and exact
  identities},'' \href{http://dx.doi.org/10.1016/j.nuclphysb.2022.116068}{{\em
  Nucl. Phys. B} {\bfseries 987} (2023) 116068},
  \href{http://arxiv.org/abs/2202.06885}{{\ttfamily arXiv:2202.06885
  [hep-th]}}.

\bibitem{Gahramanov:2013xsa}
I.~Gahramanov and G.~Vartanov, ``{Extended global symmetries for 4D $N$ = 1
  SQCD theories},''
  \href{http://dx.doi.org/10.1088/1751-8113/46/28/285403}{{\em J. Phys.}
  {\bfseries A46} (2013) 285403},
\href{http://arxiv.org/abs/1303.1443}{{\ttfamily arXiv:1303.1443 [hep-th]}}.
%%CITATION = ARXIV:1303.1443;%%.

\bibitem{Dimofte:2012pd}
T.~Dimofte and D.~Gaiotto, ``{An E7 Surprise},''
  \href{http://dx.doi.org/10.1007/JHEP10(2012)129}{{\em JHEP} {\bfseries 10}
  (2012) 129}, \href{http://arxiv.org/abs/1209.1404}{{\ttfamily arXiv:1209.1404
  [hep-th]}}.

\bibitem{Spiridonov:2008zr}
V.~Spiridonov and G.~Vartanov, ``{Superconformal indices for N = 1 theories
  with multiple duals},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2009.08.022}{{\em Nucl. Phys. B}
  {\bfseries 824} (2010) 192--216},
  \href{http://arxiv.org/abs/0811.1909}{{\ttfamily arXiv:0811.1909 [hep-th]}}.

\bibitem{Rosengren:2016mnw}
H.~Rosengren, ``{Rahman's biorthogonal functions and superconformal indices},''
  \href{http://arxiv.org/abs/1612.05051}{{\ttfamily arXiv:1612.05051
  [math.CA]}}.

\bibitem{Catak:2022glx}
E.~Catak, ``{A remark on the q-hypergeometric integral Bailey pair and the
  solution to the star-triangle equation},''
  \href{http://arxiv.org/abs/2212.13528}{{\ttfamily arXiv:2212.13528
  [math.CA]}}.

\bibitem{Schlosser:2017ixz}
M.~J. Schlosser, ``{Hypergeometric and basic hypergeometric series and
  integrals associated with root systems},''
  \href{http://arxiv.org/abs/1705.09221}{{\ttfamily arXiv:1705.09221
  [math.CA]}}.

\bibitem{Kapustin:2011jm}
A.~Kapustin and B.~Willett, ``{Generalized Superconformal Index for Three
  Dimensional Field Theories},''
\href{http://arxiv.org/abs/1106.2484}{{\ttfamily arXiv:1106.2484 [hep-th]}}.
%%CITATION = ARXIV:1106.2484;%%.

\bibitem{Gahramanov:2013rda}
I.~Gahramanov and H.~Rosengren, ``{A new pentagon identity for the tetrahedron
  index},'' \href{http://dx.doi.org/10.1007/JHEP11(2013)128}{{\em JHEP}
  {\bfseries 11} (2013) 128},
\href{http://arxiv.org/abs/1309.2195}{{\ttfamily arXiv:1309.2195 [hep-th]}}.
%%CITATION = ARXIV:1309.2195;%%.

\bibitem{Gahramanov:2014ona}
I.~Gahramanov and H.~Rosengren, ``{Integral pentagon relations for 3d
  superconformal indices},''
  \href{http://dx.doi.org/10.1090/pspum/093/01569}{{\em Proc. Symp. Pure
  Math.93,165} (2016) },
\href{http://arxiv.org/abs/1412.2926}{{\ttfamily arXiv:1412.2926 [hep-th]}}.
%%CITATION = ARXIV:1412.2926;%%.

\bibitem{Gahramanov:2015tta}
I.~Gahramanov, ``{Mathematical structures behind supersymmetric dualities},''
  \href{http://dx.doi.org/10.5817/AM2015-5-273}{{\em Archivum Math.} {\bfseries
  51} (2015) 273--286}, \href{http://arxiv.org/abs/1505.05656}{{\ttfamily
  arXiv:1505.05656 [math-ph]}}.

\bibitem{Gahramanov:2021pgu}
I.~Gahramanov and O.~E. Kaluc, ``{Bailey pairs for the q-hypergeometric
  integral pentagon identity},''
  \href{http://arxiv.org/abs/2111.14793}{{\ttfamily arXiv:2111.14793
  [math-ph]}}.

\bibitem{Bozkurt:2018xno}
D.~N. Bozkurt and I.~Gahramanov, ``{Pentagon identities arising in
  supersymmetric gauge theory computations},''
  \href{http://dx.doi.org/10.1134/S0040577919020028}{{\em Teor. Mat. Fiz.}
  {\bfseries 198} no.~2, (2019) 215--224},
  \href{http://arxiv.org/abs/1803.00855}{{\ttfamily arXiv:1803.00855
  [math-ph]}}.

\bibitem{Gahramanov:2017ysd}
I.~Gahramanov and S.~Jafarzade, ``{Integrable lattice spin models from
  supersymmetric dualities},''
  \href{http://dx.doi.org/10.1134/S1547477118060079}{{\em Phys. Part. Nucl.
  Lett.} {\bfseries 15} no.~6, (2018) 650--667},
  \href{http://arxiv.org/abs/1712.09651}{{\ttfamily arXiv:1712.09651
  [math-ph]}}.

\bibitem{Gahramanov:2017idz}
I.~Gahramanov and S.~Jafarzade, ``{Comments on the multi-spin solution to the
  Yang-Baxter equation and basic hypergeometric sum/integral identity},''
  \href{http://dx.doi.org/10.1142/S0217732319501402}{{\em Mod. Phys. Lett. A}
  {\bfseries 34} no.~18, (2019) 1950140},
  \href{http://arxiv.org/abs/1710.09106}{{\ttfamily arXiv:1710.09106
  [math-ph]}}.

\bibitem{Eren:2019ibl}
E.~Eren, I.~Gahramanov, S.~Jafarzade, and G.~Mogol, ``{Gamma function solutions
  to the star-triangle equation},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2020.115283}{{\em Nucl. Phys. B}
  {\bfseries 963} (2021) 115283},
  \href{http://arxiv.org/abs/1912.12271}{{\ttfamily arXiv:1912.12271
  [math-ph]}}.

\bibitem{Catak:2021coz}
E.~Catak, I.~Gahramanov, and M.~Mullahasanoglu, ``{Hyperbolic and trigonometric
  hypergeometric solutions to the star-star equation},''
  \href{http://dx.doi.org/10.1140/epjc/s10052-022-10661-x}{{\em Eur. Phys. J.
  C} {\bfseries 82} no.~9, (2022) 789},
  \href{http://arxiv.org/abs/2107.06880}{{\ttfamily arXiv:2107.06880
  [hep-th]}}.

\bibitem{Dimofte:2011py}
T.~Dimofte, D.~Gaiotto, and S.~Gukov, ``{3-Manifolds and 3d Indices},''
  \href{http://dx.doi.org/10.4310/ATMP.2013.v17.n5.a3}{{\em Adv. Theor. Math.
  Phys.} {\bfseries 17} no.~5, (2013) 975--1076},
  \href{http://arxiv.org/abs/1112.5179}{{\ttfamily arXiv:1112.5179 [hep-th]}}.

\bibitem{Stroganov:1979et}
{\relax Yu}.~G. Stroganov, ``{A new calculation method for partition functions
  in some lattice models},''
\href{http://dx.doi.org/10.1016/0375-9601(79)90601-7}{{\em Phys. Lett.}
  {\bfseries A74} (1979) 116--118}.
%%CITATION = PHLTA,A74,116;%%.

\bibitem{Baxter:1982xp}
R.~J. Baxter, ``{The inversion relation method for some two-dimensional exactly
  solved models in lattice statistics},''
\href{http://dx.doi.org/10.1007/BF01011621}{{\em J. Statist. Phys.} {\bfseries
  28} (1982) 1--41}.
%%CITATION = JSTPB,28,1;%%.

\bibitem{Jaekel:1982tp}
M.~T. Jaekel and J.~M. Maillard, ``{INVERSE FUNCTIONAL RELATIONS FOR LATTICE
  MODELS},'' \href{http://dx.doi.org/10.1088/0305-4470/16/9/022}{{\em J. Phys.
  A} {\bfseries 16} (1983) 1975}.

\bibitem{Perk:1986nr}
J.~H.~H. Perk and F.~Y. Wu, ``{Graphical Approach to the Nonintersecting String
  Model: Star Triangle Equation, Inversion Relation and Exact Solution},''
  \href{http://dx.doi.org/10.1016/0378-4371(86)90175-5}{{\em Physica A}
  {\bfseries 138} (1986) 100--124}.

\bibitem{Gahramanov:2022qge}
I.~Gahramanov, ``{Integrability from supersymmetric duality: a short review},''
  \href{http://arxiv.org/abs/2201.00351}{{\ttfamily arXiv:2201.00351
  [hep-th]}}.

\bibitem{Spiridonov:2010em}
V.~P. Spiridonov, ``{Elliptic beta integrals and solvable models of statistical
  mechanics},'' {\em Contemp. Math.} {\bfseries 563} (2012) 181--211,
\href{http://arxiv.org/abs/1011.3798}{{\ttfamily arXiv:1011.3798 [hep-th]}}.
%%CITATION = ARXIV:1011.3798;%%.

\bibitem{Yamazaki:2018xbx}
M.~Yamazaki, ``{Integrability As Duality: The Gauge/YBE Correspondence},''
  \href{http://dx.doi.org/10.1016/j.physrep.2020.01.006}{{\em Phys. Rept.}
  {\bfseries 859} (2020) 1--20},
  \href{http://arxiv.org/abs/1808.04374}{{\ttfamily arXiv:1808.04374
  [hep-th]}}.

\bibitem{Yagi:2015lha}
J.~Yagi, ``{Quiver gauge theories and integrable lattice models},''
  \href{http://dx.doi.org/10.1007/JHEP10(2015)065}{{\em JHEP} {\bfseries 10}
  (2015) 065},
\href{http://arxiv.org/abs/1504.04055}{{\ttfamily arXiv:1504.04055 [hep-th]}}.
%%CITATION = ARXIV:1504.04055;%%.

\bibitem{Csaki:1996sm}
C.~Csaki, M.~Schmaltz, and W.~Skiba, ``{A Systematic approach to confinement in
  N=1 supersymmetric gauge theories},''
  \href{http://dx.doi.org/10.1103/PhysRevLett.78.799}{{\em Phys. Rev. Lett.}
  {\bfseries 78} (1997) 799--802},
  \href{http://arxiv.org/abs/hep-th/9610139}{{\ttfamily arXiv:hep-th/9610139}}.

\bibitem{Csaki:1998th}
C.~Csaki, ``{The Confining N=1 supersymmetric gauge theories: A Review},'' in
  {\em {3rd Workshop on Continuous Advances in QCD (QCD 98)}}, pp.~489--499.
\newblock 4, 1998.
\newblock \href{http://arxiv.org/abs/hep-th/9807222}{{\ttfamily
  arXiv:hep-th/9807222}}.

\bibitem{Csaki:1998fm}
C.~Csaki and H.~Murayama, ``{New confining N=1 supersymmetric gauge
  theories},'' \href{http://dx.doi.org/10.1103/PhysRevD.59.065001}{{\em Phys.
  Rev. D} {\bfseries 59} (1999) 065001},
  \href{http://arxiv.org/abs/hep-th/9810014}{{\ttfamily arXiv:hep-th/9810014}}.

\bibitem{Nii:2018wwj}
K.~Nii, ``{Confinement in 3d $ \mathcal{N} $ = 2 Spin(N) gauge theories with
  vector and spinor matters},''
  \href{http://dx.doi.org/10.1007/JHEP03(2019)113}{{\em JHEP} {\bfseries 03}
  (2019) 113}, \href{http://arxiv.org/abs/1810.06618}{{\ttfamily
  arXiv:1810.06618 [hep-th]}}.

\bibitem{Nii:2019dwi}
K.~Nii, ``{Confinement in 3d $\mathcal{N}=2$ exceptional gauge theories},''
  \href{http://arxiv.org/abs/1906.10161}{{\ttfamily arXiv:1906.10161
  [hep-th]}}.

\bibitem{Kels:2013ola}
A.~P. Kels, ``{A new solution of the star-triangle relation},''
  \href{http://dx.doi.org/10.1088/1751-8113/47/5/055203}{{\em J. Phys.}
  {\bfseries A47} (2014) 055203},
\href{http://arxiv.org/abs/1302.3025}{{\ttfamily arXiv:1302.3025 [math-ph]}}.
%%CITATION = ARXIV:1302.3025;%%.

\bibitem{Kels:2017vbc}
A.~P. Kels and M.~Yamazaki, ``{Lens elliptic gamma function solution of the
  Yang\textendash{}Baxter equation at roots of unity},''
  \href{http://dx.doi.org/10.1088/1742-5468/aaa8fd}{{\em J. Stat. Mech.}
  {\bfseries 1802} no.~2, (2018) 023108},
  \href{http://arxiv.org/abs/1709.07148}{{\ttfamily arXiv:1709.07148
  [math-ph]}}.

\bibitem{Mullahasanoglu:2021xyf}
M.~Mullahasanoglu and N.~Tas, ``{Lens Partition Functions and Integrability
  Properties},'' \href{http://arxiv.org/abs/2112.15161}{{\ttfamily
  arXiv:2112.15161 [hep-th]}}.

\bibitem{de-la-Cruz-Moreno:2020xop}
J.~de-la Cruz-Moreno and H.~Garc\'\i{}a-Compe\'an, ``{Star-triangle type
  relations from $2d$ $\mathcal{N}=(0,2)$ $USp(2N)$ dualities},''
  \href{http://dx.doi.org/10.1007/JHEP01(2021)023}{{\em JHEP} {\bfseries 01}
  (2021) 023}, \href{http://arxiv.org/abs/2008.02419}{{\ttfamily
  arXiv:2008.02419 [hep-th]}}.

\bibitem{Dede:2022ofo}
M.~Dede, ``{A comment on the solutions of the generalized Faddeev-Volkov
  model},'' \href{http://arxiv.org/abs/2206.14271}{{\ttfamily arXiv:2206.14271
  [math-ph]}}.

\end{thebibliography}\endgroup

\documentclass[12pt,urlcolor=black, linkcolor=black]{article} %
\usepackage{cite}

\usepackage{amsmath, amsthm, amssymb,slashed}
\usepackage{amsfonts,amssymb, amscd,amsmath,latexsym,amsbsy,bm}
\usepackage{mathtools}
\usepackage[sort, numbers]{natbib}

\usepackage{ifpdf}
\ifpdf
  \usepackage[pdftex]{graphicx}
  \usepackage{epstopdf}
\else
  \usepackage[dvips]{graphicx}
\fi
\usepackage[letterpaper]{geometry}
\usepackage{tablefootnote}

\usepackage[usenames, dvipsnames]{xcolor}

\allowdisplaybreaks[1]

\usepackage{color}
\usepackage[colorlinks,citecolor=blue]{hyperref}

\usepackage[bbgreekl]{mathbbol}
\usepackage{amscd}

\usepackage{enumitem,cleveref}%

\usepackage{comment}
\usepackage{verbatim}

\usepackage{tikz}
\usetikzlibrary{calc}

\usepackage[all,cmtip]{xy}
\usepackage{tikz-cd}

\usetikzlibrary{matrix}
\usetikzlibrary{knots}

\usetikzlibrary{calc}

\usepackage[all,cmtip]{xy}
\usepackage{tikz-cd}

\usepackage{datetime}

\newenvironment{myfont}[2][]{\csname#2\endcsname[#1]}{}

\usepackage{makecell} 

\usepackage{slashed}
\usepackage[makeroom]{cancel}
\usepackage[normalem]{ulem} 
\usepackage{soul}

\begin{document}
\begin{titlepage}
\begin{flushright}
\end{flushright}
\vskip 12.mm

\begin{center}

{\bf\Large{
Comments on flavor symmetry breaking and three-dimensional superconformal index
\\[6.8mm] 
}}

\vskip.5cm
\quad\quad\quad
\large{ \textbf{Ilmar Gahramanov} }
\hfill  


\vskip.5cm
{\small{\textit{$^1$ {Department of Physics, Bogazici University, 34342 Bebek, Istanbul, Turkey} \\}}
}
\vskip.2cm
{\small{\textit{$^2$Steklov Mathematical Institute of Russ. Acad. Sci.,
Gubkina str. 8, 119991 Moscow, Russia}\\}}
\vskip.2cm
{\small{\textit{$^3$ Department of Mathematics, Khazar University,  Mehseti St. 41, AZ1096, Baku, Azerbaijan}\\}}
\vskip.2cm
{\small{\textit{$^4$ Institute of Radiation Problems ANAS, B.Vahabzade 9, AZ1143 Baku, Azerbaijan}\\}}

\end{center}
\vskip.35cm
\baselineskip 12pt
\begin{abstract}

We investigate the superconformal index of a three-dimensional $SU(2)$ ${\mathcal N}=2$ supersymmetric QCD  with $N_f=4$ flavors. This theory confines with the breaking of a global symmetry. We obtain the index of this theory by integrating out a flavor from the theory with $N_f=6$ flavors. The superconformal index of the theory vanishes for generic values of the flavor fugacities. However, the specific choice of fugacities allows us to describe the flavor symmetry-breaking phenomenon.

 


\end{abstract}

\end{titlepage}



  \pagenumbering{arabic}
    \setcounter{page}{2}
    



\noindent

\section{Introduction}
	
Supersymmetric duality relates the infrared limits of different supersymmetric gauge theories. The basic example of such duality \cite{Seiberg:1994pq} is a four-dimensional $SU(N_c)$ ``electric'' gauge theory with $N_f$ flavors of quarks in fundamental and antifundamental representations which possesses a dual description in terms of $N_f$ ``magnetic'' flavors of quarks charged under $SU(N_f-N_c)$ gauge group. For $N_f = N_c$, the infrared theory is confined, and the classical constraint on the meson and baryon fields is modified by one instanton effect, leading to chiral symmetry breaking. In case $N_c=2$ the global symmetry $SU(4)$ is broken\footnote{The same happens for the two colors non-supersymmetric QCD with the two flavors which enjoys the extended chiral symmetry $SU(4)$ instead of $SU(2)\times SU(2)$, see e.g. \cite{Smilga:1994tb,Kogut:1999iv}, see, e.g. \cite{Magnea:1999iv,Dunne:2002vb} for the three-dimensional case.} to its maximal vector-like subgroup $SP(4)$ by the chiral condensate \cite{Seiberg:1994bz}. The chiral symmetry breaking for four-dimensional ${\mathcal N}=1$ supersymmetric gauge theories was considered on the level of the superconformal indices in \cite{Spiridonov:2014cxa} by Spiridonov and Vartanov.

Three-dimensional $\mathcal N=2$ supersymmetric gauge theories have rich structures \cite{Intriligator:1996ex,deBoer:1996mp,Aharony:1997bx,deBoer:1997kr,Amariti:2016kat}, such as duality, confinement, and chiral symmetry breaking, similar to four-dimensional $\mathcal N=1$ supersymmetric gauge theories\footnote{The three-dimensional $\mathcal N=2$ theory has the same number of supercharges as a four-dimensional $\mathcal N=1$ theory. In fact, the three-dimensional $\mathcal N=2$ theories discussed in this work are the dimensional reduction of the corresponding four-dimensional $\mathcal N=1$ theories \cite{Dolan:2011rp,Spiridonov:2014cxa, Aharony:2013dha, Amariti:2015xna}.}. In this work, we consider the flavor symmetry-breaking phenomenon for the three-dimensional $SU(2)$ $\mathcal N=2$ theory with $N_f=4$ flavors by using the superconformal index technique. Similar to the case of the three-dimensional $S^3$ sphere partition function case \cite{Spiridonov:2014cxa} the superconformal index of such a theory involves the Dirac delta functions which reflect the presence of symmetry breaking
\begin{align} \nonumber
		\sum_{m \in {\mathbb Z}} \oint \frac{(q^{1+(n_\alpha\pm n_w\pm m)/2} \alpha^{-1} w^{\mp} z^{\mp})_\infty}{(q^{n_\alpha \pm n_w \pm m} \alpha w^\pm z^\pm)_\infty}  \frac{(q^{1+(-n_\alpha\pm n_y\pm m)/2} \alpha y^{\mp} z^{\mp})_\infty}{(q^{-n_\alpha \pm n_y \pm m} \alpha^- y^\pm z^\pm)_\infty} \frac{(1-q^m z^{\pm 2})}{q^m} \frac{dz}{4\pi i z} \\ \label{invrel}
		= \frac{(q^{1+n_\alpha} \alpha^{\mp 2};q)_\infty}{(q^{n_\alpha} \alpha^{\pm 2};q)_\infty} \frac{(q^{1 \pm n_y} w^{ \mp 2};q)_\infty}{(q^{\pm n_w} w^{\pm 2};q)_\infty} \left(\delta(\theta+\chi) \delta_{n_w+n_y,0}+\delta(\theta-\chi)\delta_{n_w-n_y,0}\right) \;,
\end{align}
where we use the usual $q$-Pochhammer symbol $(z;q)_\infty=\prod_{i=0}^{\infty} (1-zq^i)$ and shorthand notations $f(z^\pm)=f(z) f(\frac{1}{z})$. On the left-hand side of equality (\ref{invrel})	we have the three-dimensional $\mathcal N=2$ superconformal index of a theory with $SU(2)$ gauge group and $N_f=4$ chiral fields with the naive flavor group $SU(2)\times SU(2)$.
However, the true flavor group is $(SU(2)\times SU(2))_{diag}$ and the index has non-zero support only on the corresponding
subset of flavor fugacities. This is precisely the manifestation of chiral symmetry breaking in confining theories similar to the three-dimensional squashed-sphere partition functions case \cite{Spiridonov:2014cxa}. Note that the expression (\ref{invrel}) plays a role of the local inversion relation in integrable lattice models \cite{Gahramanov:2015cva,Gahramanov:2016ilb,Kels:2015bda}.

One can consider the derivation of (\ref{invrel}) as another algebraic approach to the inversion relation. 


	\subsection{Superconformal index}

We consider supersymmetric index of a three--dimensional ${\mathcal N}=2$ superconformal field theory\footnote{For more detailed introduction, see e.g. \cite{Krattenthaler:2011da,Gahramanov:2015cva,Gahramanov:2016wxi}.} which can be defined as a partition function on $S^2 \times S^1$ with the periodic boundary condition of the fields along $S^1$ \cite{Bhattacharya:2008zy,Bhattacharya:2008bja,Kim:2009wb,Imamura:2011su}
	\begin{equation}
		{I}(q,\{ a_i \})=\text{Tr} \left[ (-1)^\text{F} e^{-\beta\{Q, Q^\dagger \}} q^{\frac12 (\Delta+j_3)}\prod_i
		a_i^{F_i} \right] \;,
	\end{equation}
where the trace is taken over the Hilbert space of the theory on $S^2$. The $Q$ is a certain supersymmetric charge in three-dimensional ${\mathcal N}=2$ superconformal algebra which satisfies the anti-commutation relation $\{Q, Q^{\dagger}\}=\Delta-R-j_3$, where $\Delta$ is the energy, $j_3$ is the third component of the angular momentum on $S^2$, $R$ is the R-charge. The operator $\text{F}$ plays a role of the fermion number, $F_i$ is the charge of global symmetry commuting with supercharge $Q$, and $a_i$ is the corresponding fugacity. 

The superconformal index can be computed exactly by the supersymmetric localization technique \cite{Willett:2016adv} and it takes the form of the following matrix integral\footnote{We skip the contribution of the Chern--Simons term, since in this work we consider theories without this term.} 
	\begin{equation}
		I(q, \{a_j\}, \{n_j\}) = \sum_{m_i}\frac{1}{|W_m|} \oint \prod_{i=1}^{\text{rank}G} \frac{dz_i}{2 \pi i z_i} Z_{\text{gauge}} (z_i,m_i; q) Z_{\text{chiral}}(z_i,m_i; a_j, n_j; q) \;.
	\end{equation}
Here $Z_{\text{gauge}}$ and $Z_{\text{chiral}}$ stand for the contributions from the vector and chiral multiplets, respectively, $|W_m|$ is the dimension of the Weyl group, the sum is over monopole configurations, and the integral is over gauge fugacities. The form of the index is completely determined if one knows group-theoretical data, i.e. symmetries and corresponding representations of a theory.  The reader interested in the precise formulation of the three-dimensional superconformal index can refer to the papers \cite{Imamura:2011su,Gahramanov:2016wxi}.

\subsection{Duality for 3d $SU(2)$  $\mathcal N = 2$ SQCD with $N_f = 6$ flavor}

As an example, we will be studying the global symmetry breaking in three-dimensional $\mathcal{N }=2$ supersymmetric QCD with $SU(2)$ gauge group and $N_f=4$ flavors of quarks. 	To obtain such theory one can take first the $s$-confining theory\footnote{Three-dimensional s-confining theories were discussed in several works, e.g. \cite{Csaki:2014cwa,Nii:2018erm,Nii:2019ebv,Amariti:2022wae} and the superconformal indices for such theories were presented in \cite{Gahramanov:2013xsa,Gahramanov:2016wxi}.} with the gauge group $SU(2)$ and $N_f=6$ flavors. In the case of $SU(2)$ gauge group quarks and antiquarks transform in the same way under the action of the gauge group, i.e. the fundamental and antifundamental representations are equivalent, and therefore the flavor group enhanced to $SU(6)$ symmetry. The superconformal index of the theory can be expressed in the following form
	\begin{equation} \label{electric}
		I_{E} \ = \ \sum_{m=-\infty}^{\infty}  \oint \prod_{i=1}^6 \frac{(q^{1+(m+n_i)/2}/a_iz,q^{1+(n_i-m)/2} z/a_i ;q)_\infty}{(q^{(m+n_i)/2}a_iz,q^{(n_i-m)/2} a_i/z ;q)_\infty} \frac{(1-q^m z^2)(1-q^m z^{-2})}{q^m z^{6m}} \frac{dz}{4\pi i z}
	\end{equation} 
	with the balancing conditions
	\begin{equation} \label{balcon}
		\prod_{i=1}^6 a_i = q \;\; \text{and} \;\; \sum_{i=1}^6 n_i =  0 \;.
	\end{equation}
The variables $a_i$ and $z$ are the fugacities associated with the $SU(6)$ flavor group and $SU(2)$ gauge group, respectively. In \cite{Dimofte:2012pd}, it was shown that the superconformal index of this theory has an extended $SO(12)$ flavor symmetry when coupled to four-dimensional multiplets with specific boundary conditions.

It is known that there is magnetic dual theory without gauge degrees of freedom and with fifteen chiral multiplets in the totally antisymmetric tensor representation of the same flavor group. The superconformal index of the magnetic theory is
	\begin{equation}
		I_M  = \frac {1}{ \prod_{i=1}^6 a_i^{n_i}}  \prod_{1 \leq i<j  \leq 6} \frac {(q^{1+(n_i+n_j)/2}/a_i a_j;q)_\infty}{(q^{(n_i+n_j)/2}a_i a_j;q)_\infty} \;.
	\end{equation}

Since these theories are IR dual, the superconformal indices need to be in agreement \cite{Spiridonov:2008zr, Gahramanov:2016wxi}. The result is given by the following q-hypergeometric identity
	\begin{align} \nonumber
		&  \sum_{m=-\infty}^{\infty}  \oint \prod_{i=1}^6 \frac{(q^{1+(m+n_i)/2}/a_iz,q^{1+(n_i-m)/2} z/a_i ;q)_\infty}{(q^{(m+n_i)/2}a_iz,q^{(n_i-m)/2} a_i/z ;q)_\infty} \frac{(1-q^m z^{\pm 2})}{q^m z^{6m}} \frac{dz}{4\pi i z} \\ \label{mainint}
		& \qquad \qquad \qquad \qquad =  \frac {1}{ \prod_{i=1}^6 a_i^{n_i}}  \prod_{1 \leq i<j  \leq 6} \frac {(q^{1+(n_i+n_j)/2}/a_i a_j;q)_\infty}{(q^{(n_i+n_j)/2}a_i a_j;q)_\infty}  
	\end{align}
with the balancing conditions (\ref{balcon}). This integral identity was proven in \cite{Gahramanov:2016wxi} and  studied in \cite{Gahramanov:2016wxi,Gahramanov:2015cva,Rosengren:2016mnw}. 

The identity (\ref{mainint}) can be written as a star-triangle relation (the non-operator form of the Yang-Baxter equation) \cite{Gahramanov:2015cva,Kels:2015bda} which is a sufficient condition for integrability for the Ising-type lattice spin models. The Bailey pairs  and the operator form of the Yang-Baxter equation related to this integral identity also were constructed in \cite{Gahramanov:2015cva} (see, also \cite{Catak:2022glx}). 

Note that the basic hypergeometric integral identity (\ref{mainint}) is new in the literature and needs detailed study and classification \cite{Schlosser:2017ixz}. Such identities (containing summation and integration) recently appear in supersymmetric gauge theory computations \cite{Krattenthaler:2011da,Kapustin:2011jm,Aharony:2013dha,Gahramanov:2013rda,Gahramanov:2014ona, Gahramanov:2015tta,Gahramanov:2016wxi,Rosengren:2016mnw,Gahramanov:2021pgu,Bozkurt:2018xno} and integrable models of statistical mechanics \cite{Gahramanov:2015cva,Gahramanov:2017ysd,Gahramanov:2017idz,Eren:2019ibl, Catak:2021coz, Catak:2022glx}.

\section{Breaking of flavor symmetry}
	
Our goal will be to reduce the number of flavors from $N_f=6$ to $N_f=4$ case for the integral identity (\ref{mainint}).  One can perform the reduction of superconformal indices (\ref{mainint}) by taking the following limit\footnote{Note that in work \cite{Gahramanov:2013xsa}, the authors take another limit $a_5 a_6=q^{1/2}$ and obtain the s-confining theory without breaking $SU(4)$ flavor symmetry. In \cite{Gahramanov:2015cva}, authors considered superconformal indices of dual theories with $SU(4)$ flavor symmetry without superpotential term (without balancing condition).}
	\begin{equation} 
		a_5 a_6  =  q, \;\; n_5+n_6=0
	\end{equation}
Generically this limit gives zero on both sides, however for special choices of fugacities $a_i$ one can deal with such integrals. Let us denote
	\begin{equation}\label{pars1}
		a_1=\alpha w, \quad a_2=\alpha w^{-1},\quad
		a_3=\beta y,\quad  a_4=\beta y^{-1},
	\end{equation}
	and the corresponding discrete parameters
	\begin{equation}\label{pars2}
		n_1=n_\alpha+n_w, \quad n_2=n_\alpha- n_{w},\quad
		n_3=n_\beta+n_y,\quad  n_4=n_\beta-n_{y}.
	\end{equation}
	In terms of new variables, the balancing conditions read as
	\begin{equation}
		a_5 a_6 \alpha^2 \beta^2 = q \;, \;\;\; 2n_\alpha+ 2 n_\beta+n_5+n_6=0 \;.
	\end{equation}
To proceed further we need to consider the following integral: multiply the integral (\ref{electric})  by a holomorphic function near the unit circle $f(y)$ and integrate over the variable  $y$. Then using the integral identity (\ref{mainint}) we obtain the following expression\footnote{Here we assume that $\alpha^2 \neq 1$.}
	\begin{align} \nonumber
		\int_{\mathbb{T}} &
		f(y)  I_E \frac{dy}{2 \pi \textup{i} y}
		 = \alpha^{-2n_\alpha} w^{-2n_w} \beta^{-2n_\beta}  a_5^{-n_5} a_6^{-n_6} \frac{(q^{1+n_\alpha} \alpha^{-2};q)_\infty}{(q^{n_\alpha} \alpha^2;q)_\infty} \frac{(q^{1+n_\beta} \beta^{-2};q)_\infty}{(q^{n_\beta} \beta^2;q)_\infty} \\ \nonumber
		& \quad \times  \frac{(q^{1+\frac{n_5+n_6}{2}} (a_5 a_6)^{-1};q)_\infty}{(q^{\frac{n_5+n_6}{2}} a_5 a_6;q)_\infty}
		 \frac{(q^{1+\frac{n_\alpha \pm n_w+n_5}{2}} (\alpha w^{\pm 1} a_5)^{-1};q)_\infty}{(q^{\frac{n_\alpha \pm n_w+n_5}{2}} \alpha w^{\pm 1} a_5;q)_\infty}
		\frac{(q^{1+\frac{n_\alpha \pm n_w+n_6}{2}} (\alpha w^{\pm 1} a_6)^{-1};q)_\infty}{(q^{\frac{n_\alpha \pm n_w+n_6}{2}} \alpha w^{\pm 1} a_6;q)_\infty} \\ 
		& \quad \times
		\int_{\mathbb{T}} \frac{dy}{2 \pi \textup{i} y} f(y) \; y^{-2n_y} \; \frac{(q^{1+\frac{n_\alpha+n_\beta \pm n_w \pm n_y}{2}} (\alpha \beta w^{\pm 1} y^{\pm 1})^{-1};q)_\infty}{(q^{\frac{n_\alpha+ n_\beta \pm n_w \pm n_y}{2}} \alpha \beta w^{\pm 1} y^{\pm 1};q)_\infty}   \prod_{j=5}^{6} \frac{(q^{1+\frac{n_\beta \pm n_y+n_j}{2}} (\beta y^{\pm 1}  a_j)^{-1};q)_\infty}{(q^{\frac{n_\beta \pm n_y+n_j}{2}} \beta y^{\pm 1} a_j;q)_\infty} \;. \label{fyint}
	\end{align}
	The first term in the integrand of (\ref{fyint}) has the following sequences of poles
	\begin{align}
		y_{in} & = \alpha \beta
		w^{\pm1} q^{\frac{n_\alpha+n_\beta \pm n_w-n_y}{2}} q^{j}, \\
		y_{out} & = \frac{1}{\alpha \beta} w^{\pm1} q^{\frac{-n_\alpha-n_\beta \pm n_w-n_y}{2}} q^{-j},
	\end{align}
	where $i,j \in \mathbb{Z}_{\geq 0}$ and the {\it in}-poles converge to zero $y=0$ and {\it out}-poles go to infinity.  In order to calculate the integral (\ref{fyint}) we deform the counter $\mathbb T$ such that only the residues of the in-poles are picked up and there are no singularities lying on the counter $C$. Then  the residue calculation leads  to the result
\begin{align} \nonumber
		\int_{\mathbb{T}}
		& f(y)I_E \frac{dy}{2 \pi \textup{i} y}
		=\alpha^{-2n_\alpha} w^{-2n_w} \beta^{-2n_\beta}  a_5^{-n_5} a_6^{-n_6} \frac{(q^{1+n_\alpha} \alpha^{-2};q)_\infty}{(q^{n_\alpha} \alpha^2;q)_\infty} \frac{(q^{1+n_\beta} \beta^{-2};q)_\infty}{(q^{n_\beta} \beta^2;q)_\infty} \\ 
		& \times \frac{(q^{1+\frac{n_5+n_6}{2}} (a_5 a_6)^{-1};q)_\infty}{(q^{\frac{n_5+n_6}{2}} a_5 a_6;q)_\infty} \prod_{j=5}^{6} \frac{(q^{1+\frac{n_\alpha \pm n_w+n_j}{2}} (\alpha w^{\pm 1} a_j)^{-1};q)_\infty}{(q^{\frac{n_\alpha \pm n_w+n_j}{2}} \alpha w^{\pm 1} a_j;q)_\infty}  \left( I(w) + I(w^{-1}) +I(C) \right)
\end{align}
where
\begin{align} \nonumber
 I(w) & =    f\left(\alpha\beta w q^{\frac{n_\alpha+n_\beta+n_w-n_y}{2}}\right) \; (\alpha\beta w q^{\frac{n_\alpha+n_\beta+n_w-n_y}{2}})^{-2n_y} \;
	 \frac{(q^{1+ny}  (\alpha \beta w)^{-2};q)_\infty} 
		{(q^{n_\alpha+ n_\beta + n_w } (\alpha \beta w)^2;q)_\infty}\\ \nonumber
  	& \quad \times	\frac{(q^{1+n_\alpha+n_\beta+n_w-n_y};q)_\infty} 
		{(q;q)_\infty}
		\frac{(q^{1-n_w+n_y}  (\alpha \beta)^{-2};q)_\infty} 
		{(q^{n_\alpha+ n_\beta } (\alpha \beta)^2;q)_\infty}
		\frac{(q^{1+n_\alpha+n_\beta-n_y}  w^2;q)_\infty} 
		{(q^{- n_w}  w^{-2};q)_\infty} \\
	& \quad \times	\prod_{i=5}^6
		\frac{(q^{1+\frac{-n_\alpha+2n_y+n_i}{2}}  (\alpha \beta^2 w a_i)^{-1};q)_\infty} 
		{(q^{\frac{2n_\beta+ n_\alpha+n_w+n_i}{2} } \alpha \beta^2 w a_i ;q)_\infty}
		\frac{(q^{1+\frac{2n_\beta+n_\alpha+n_i+n_w-2n_y}{2} }  ( \alpha w^{-1} a_i)^{-1};q)_\infty} 
		{(q^{\frac{- n_w-n_\alpha+n_i}{2}}   \alpha w^{-1} a_i;q)_\infty}
\end{align}
\begin{align} \nonumber
 I(w^{-1}) & =  f\left(\alpha\beta w^{-1} q^{\frac{n_\alpha+n_\beta-n_w-n_y}{2}}\right) (\alpha\beta w^{-1} q^{\frac{n_\alpha+n_\beta-n_w-n_y}{2}})^{-2n_y}  \frac{(q^{1+n_w+n_y}  (\alpha \beta)^{-2};q)_\infty} 		{(q^{n_\alpha+ n_\beta} (\alpha \beta)^2;q)_\infty}\\ \nonumber
		& \quad \times \frac{(q^{1+n_\alpha+n_\beta-n_y} w^{-2};q)_\infty} 
		{(q^{n_w} w^2;q)_\infty} 
		\frac{(q^{1+n_y}  (\alpha \beta w^{-1})^{-2};q)_\infty} 
		{(q^{n_\alpha+ n_\beta - n_w} (\alpha \beta w^{-1})^2;q)_\infty}
		\frac{(q^{1+n_\alpha+n_\beta-n_w-n_y};q)_\infty} 
		{(q;q)_\infty} \\ 
		& \quad \times 		\prod_{i=5}^6
		\frac{(q^{1+\frac{-n_\alpha+2n_y+n_i}{2}}  (\alpha \beta^2 w^{-1} a_i)^{-1};q)_\infty} 
		{(q^{\frac{2n_\beta+ n_\alpha-n_w+n_i}{2} } \alpha \beta^2 w^{-1} a_i ;q)_\infty}
		\frac{(q^{1+\frac{2n_\beta+n_\alpha+n_i-n_w-2n_y}{2} }  ( \alpha w a_i)^{-1};q)_\infty} 
		{(q^{\frac{n_w-n_\alpha+n_i}{2}}   \alpha w a_i;q)_\infty}
  \end{align}
\begin{align} 
 I(C) & =  \int_{\mathcal C}  \frac{dy}{2 \pi \textup{i} y} f(y) y^{-2n_y} \frac{(q^{1+\frac{n_\alpha+n_\beta \pm n_w \pm n_y}{2}} (\alpha \beta w^{\pm 1} y^{\pm 1})^{-1};q)_\infty}{(q^{\frac{n_\alpha+ n_\beta \pm n_w \pm n_y}{2}} \alpha \beta w^{\pm 1} y^{\pm 1};q)_\infty} 
		\prod_{i=5}^6 \frac{(q^{1+\frac{n_\beta \pm n_y+n_i}{2}} (\beta y^{\pm 1} a_i)^{-1};q)_\infty}{(q^{\frac{n_\beta \pm n_y+n_i}{2}} \beta y^{\pm 1} a_i;q)_\infty} \label{intC}
\end{align}

In the limit  $\alpha\beta\to 1$ and $n_\alpha+n_\beta \to 0$, the poles at $ \alpha \beta w^{\pm1} q^{\pm n_w-n_y}$ and
	$w^{\pm1} q^{\pm n_w-n_y}/{\alpha \beta}$ start to pinch $\mathbb{T}$. The integral (\ref{intC}) vanishes in this limit. As a result, we obtain\footnote{In order to cancel out many terms we used the  equality \cite{Dimofte:2011py,Gahramanov:2015cva}
		\begin{equation}
			\frac{ (q^{1-m/2}z^{-1} ;q)_\infty} {( q^{-m/2} z;q)_\infty }
			=\frac{q^{m/2}}{(-z)^m} \frac{ (q^{1+m/2}z^{-1} ;q)_\infty} {( q^{+m/2} z;q)_\infty }, \quad m\in {\mathbb Z}.
			\label{identity}\end{equation}.}
	\begin{align} \nonumber
		\int_{\mathbb{T}}
		f(y)I_E \frac{dy}{2 \pi \textup{i} y}
		& = \frac{(q^{1+n_\alpha} \alpha^{\mp 2};q)_\infty}{(q^{n_\alpha} \alpha^{\pm 2};q)_\infty} \frac{(q^{1 \pm n_y} w^{ \mp 2};q)_\infty}{(q^{\pm n_w} w^{\pm 2};q)_\infty} \\ \label{prev}
		& \quad \times \left( f(w q^{(n_w-n_y)/2}) \delta_{n_w+n_y,0}
		+ f(w^{-1} q^{(-n_w-n_y)/2}) \delta_{n_w-n_y,0} \right)
	\end{align}
Here $\delta_{i,j}$ is the Kronecker delta. Now we	denote $y=e^{2\pi\textup{i} \theta}, w=e^{2\pi\textup{i} \chi}$
	and pass to the integration over real variable $\theta$.
	Because of the arbitrariness of the function $f(y)$, we can
	give to the function $I_E$ a distributional sense and rewrite the expression (\ref{prev}) in the following form
	\begin{align} \nonumber
		\sum_{m \in {\mathbb Z}} \oint \frac{(q^{1+(n_\alpha\pm n_w\pm m)/2} \alpha^{-1} w^{\mp} z^{\mp})_\infty}{(q^{n_\alpha \pm n_w \pm m} \alpha w^\pm z^\pm)_\infty}  \frac{(q^{1+(-n_\alpha\pm n_y\pm m)/2} \alpha y^{\mp} z^{\mp})_\infty}{(q^{-n_\alpha \pm n_y \pm m} \alpha^- y^\pm z^\pm)_\infty} \frac{(1-q^m z^2)(1-q^m z^{-2})}{q^m} \frac{dz}{4\pi i z} \\ \label{mainresult}
		= \frac{(q^{1+n_\alpha} \alpha^{\mp 2};q)_\infty}{(q^{n_\alpha} \alpha^{\pm 2};q)_\infty} \frac{(q^{1 \pm n_y} w^{ \mp 2};q)_\infty}{(q^{\pm n_w} w^{\pm 2};q)_\infty} \left(\delta(\theta+\chi) \delta_{n_w+n_y,0}+\delta(\theta-\chi)\delta_{n_w-n_y,0}\right)
	\end{align}
	where $\delta(\theta)$ is the periodic Dirac delta function, $\delta(\phi+1)=\delta(\phi)$. The appearance of the Dirac delta functions describes the symmetry breaking of the flavor group  to the $\left( SU(2) \times SU(2) \right)_{diag}$. 
 
 The equality (\ref{mainresult}) has an interpretation as the inversion relation\footnote{Inversion relation is a simple relation as a star-triangle relation within the framework of integrable lattice spin models of statistical mechanics, see, e.g. \cite{Stroganov:1979et,Baxter:1982xp,Jaekel:1982tp,Perk:1986nr}. In the context of S-matrix for particle physics in $1+1$  dimensions it has the meaning of the unitarity relation.} for exactly solvable lattice spin model introduced in \cite{Gahramanov:2015cva,Kels:2015bda}. In particular, this observation can be used \cite{Gahramanov:2022qge} for a better understanding of the correspondence \cite{Spiridonov:2010em,Gahramanov:2017ysd,Yamazaki:2018xbx,Yagi:2015lha} between supersymmetric gauge theories and integrable models of statistical mechanics.

	\section{Concluding remarks}
 
We have shown that it is possible to obtain a three-dimensional $\mathcal N=2$ s-confining theory with flavor symmetry breaking by analyzing the superconformal index of the theory.  There are several potential directions to extend our work. It would be interesting to obtain the superconformal index with delta functions via localization technique if it is possible. 


Though studied only $SU(2)$ supersymmetric gauge theory with $N_f=4$ flavors the approach discussed here also applies to gauge theories with other flavor groups. We would expect analogous behavior of the superconformal index for the general $SU(N)$ theory, however, it is computationally complicated to work with the multivariate analog of the basic hypergeometric integral identity (\ref{mainint}).
	
The four-dimensional s-confining theories are classified \cite{Csaki:1996sm,Csaki:1998th,Csaki:1998fm}. The chiral symmetry is always broken in such theories if one reduces flavor symmetries to $N_f = N_c$. The corresponding three-dimensional analysis \cite{Csaki:2014cwa,Nii:2018wwj,Nii:2019ebv,Nii:2019dwi} for such theories need to be performed.  

Finally, there are many integrable lattice models \cite{Kels:2013ola,Gahramanov:2016ilb,Kels:2017vbc,Mullahasanoglu:2021xyf,de-la-Cruz-Moreno:2020xop,Dede:2022ofo} obtained recently via exact results from supersymmetric gauge theories. A similar approach can be used for obtaining corresponding inversion relations. We hope to elaborate on this as well as on other topics elsewhere.
	
\section*{Acknowledgements} The work originated from stimulating discussions with Vyacheslav Spiridonov, I thank him for his help, suggestions, and contributions to the
ideas presented here. I would like to thank Keita Nii, Mustafa Mullahasanoglu, Erdal Catak, Yasin Şale, Merve Kaftan, and Yağız Eren Kiremit for the discussions. This work has been supported by the Russian Science Foundation grant number 22-72-10122 2 (https://rscf.ru/en/project/22-72-10122).  I am grateful to the Institut des Hautes Etudes Scientifiques, IHES (Bures-sur-Yvette, France), for the warm hospitality, where part of this work was carried out. The visit to IHES was supported by Nokia Fund in 2023.


\bibliographystyle{utphys.bst}
\bibliography{references.bib}

\end{document}